answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability of getting a even number in a deck of cards?

Well, in a deck of cards, there's 2, 4, 6, 8, and 10, so that's five cards. Take those five cards and multiply it by four, because there's four different suits (clubs, diamonds, spades, and hearts.) This gives you 20, out of 52 cards, or a 20/52 chance. You can reduce this by dividing by 4, to get a 4/13 chance. You can use a calculator to divide that out and find the percentage answer, which is 38.5 percent.

If you include the two jokers, then you have a 20/54 chance of drawing an even number, which works out to a 37 percent chance.

Which best describes probability?

the likelihood that something will happen

What is the probability in picking a vowel in mathematical?

Probablity of picking up 3 vowels from the word MATHEMATICS Each of the letters in the word MATHEMATICS is on a letter tile in a bag

A particular fruit's weights are normally distributed, with a mean of 298 grams and a standard deviation of 13 grams.If you pick one fruit at random, what is the probability that it will weigh between 262 grams and 269 grams?

A particular fruit's weights are normally distributed, with a mean of 760 grams and a standard deviation of 15 grams. If you pick one fruit at random, what is the probability that it will weigh between 722 grams and 746 grams-----A particular fruit's weights are normally distributed, with a mean of 567 grams and a standard deviation of 25 grams.

What is the probability of picking a card with a vowel on it from a box of cards in which each letter of alphabet is written on a card?

Think, if there are A, E, I, O, U, and couting Y, there are 6 vowels. there are 26 letters. So the answer is 6 over 26 and then you simplify it to 3 over 13.

How many different ways can you arrange the letters in the word journalism?

There are 10 letters is the word JOURNALISM. Since they are all different, the number of ways you can arrange them is simply the number of permutations of 10 things taken 10 at a time, or 10 factorial, or 3,628,800.

A bag contains 3 blue marbles 7 red marbles and 5 yellow marbles. If Joe picks 1 marble from the bag at random what is the probability that it will be blue?

  1. A bag of marbles contains 13 marbles. 5 Blue, 3 Yellow, 4 Green and 1 Red. Leave all answers as a ratio in lowest terms. 18 points

On a single draw, what is the probability of drawing a yellow marble?

What is the probability of not drawing a yellow marble?

What are the odds in favor of drawing a blue marble?

What is the probability of drawing a red or yellow marble?

What is the probability of drawing a purple marble?

If you had to bet on drawing a marble of a certain color what color would you not choose?

A bag contains 3 red marbles 5 blue marbles 8 yellow marbles and 4 black marbles what is the theoretical probability of randomly picking each color marble?

The theoretical probability of randomly picking each color marble is the number of color marbles you have for each color, divided by the total number of marbles. For example, the probability of selecting a red marble is 3/20.

What is the link between the characteristic function of a discrete probability distribution and its discrete Fourier transform?

Provided the random variable can take only finitely many, equally spaced values, you can write one in terms of the other. For details see the paper by Ales Cerny entitled "Introduction to Fast Fourier Transform in finance".

Diana has a box containing 3 red and 2 green marbles. She selects 2 marbles at random. Find the probability if she replaces the 1st marble before she draws the second?

With replacement, the odds of the second extraction do not change after the first, because of independence of the two extractions. Hence, the probability of extracting two reds is (3/5)*(3/5), the probability of two greens is (2/5)*(2/5) and that of one red and one green is 2*(3/5)*(2/5).

What is the probability that the sequence will be Heads-Tails-Heads if you toss a coin 3 times?

Since there are 2 outcomes for a coin toss, and you will toss the coin 3 times the number of outcomes are 23 or 8. Since H-T-H can occur only 1 way, the probability of the H-T-H sequence is 1/8.

Why is strata called homogenous in stratified sampling?

Homogeneous refers to groups composed of parts or elements that are all of the same kind or nature. In stratified sampling, a population which is composed of diverse groupings is subdivided into two or more groups so that the diversity is decreased in the subgroups. For example, if the total population is composed of males and females, then stratification into subgroups of male and female will result in strata that are of the same kind with respect to the classification variable gender: i.e, the strata are homogeneous. Other classification variables or combinations of classification variables may be used to improve homogeneity.

Why is the mean the standard partner of the standard deviation?

The mean and standard deviation often go together because they both describe different but complementary things about a distribution of data. The mean can tell you where the center of the distribution is and the standard deviation can tell you how much the data is spread around the mean.

If you have 10 football matches with 3 outcomes win lose and draw how many different combinations are there?

IF YOU HAVE ONE GAME, YOU HAVE A POSIBILITY OF 3 OUTCOMES, IF YOU HAVE 2 GAMES IT GO'S UP TO 9 POSIBILE OUTCOMES.

FIRST ANSWER 1X3=3

SECOND IS 3X3=9

FOR 10 GAMES IT IS 3X3 TEN TIMES

3X3X3X3X3X3X3X3X3X3=59,049

POSSIBLE COMBINATIONS

(I THINK)HA HA

If a student guesses on all 5 questions on a tru-false exam what is the probability that he or she get them all right?

For each question the student has a 50% chance of answering right.

We can express this as a fraction of 1 (1 being 100%): 1/2, i.e. 0.5

There are 5 questions, so the answer will be:

0.5*0.5*0.5*0.5*0.5=0.5^5=0.03125

or 3.125%

Trending Questions
An experiment consists of rolling two fair dice and adding the dots on the two sides facing up therefore what is the probability that the sum of the dots is divisible by 2 and 3? What is a Probability is used to estimate probabilities by making certain assumptions about an experiment? In a bow with 48 cards. How many cards were there in 7 boxes? What is the primary purpose of the mean absolute deviation in forcasting? If the odds in favor is one to fifteen for winning a door prize then what is the probability of winning a door prize? Is the something wrong when you hear your heatbeat pulsate in your eardrum? How many red jacks are there in a set of cards? What is another name for answer? How many outcomes are possible if each coin is flipped once? What is the probability of obtaining exactly three heads in four flips of a coin given that at least two are heads? Given the mean of 259 and a standard deviation of 229.1 what is the z-score of the value 78? The probability that Bob will win a game of snooker is 0.7 and the probability that he will win a game of chess is 0.6 What is the probability that he will not win both games? A nursing class has 13 women and 18 men If a student is chosen randomly to be the team leader what is the probability the student is a woman? What is the event space for rolling a number greater than 2 with one die? How many ways can the letters in glasses be arranged? If you draw a card from a normal deck of cards what is the probability the card is an jack of diamonds? What are the odds in favor of rolling a 2 or 3 on a single roll of a fair die? How do you do probability saying no? Taylor tossed one green number cube and one red number cube each cube is numbered from 1-6what is the probility that the sum of the numbers on the cube is 4 write the answer as a fraction in simplest? The likelihood that an event will occur like tossing a coin?